An aluminum cylinder is 10.0 cm in length and has a radius of 0.25 cm. If the mass of a single Al atom is 4.48 × 10−23g, calculate the number of Al atoms present in the cylinder. The density of aluminum is 2.70 g/cm3 .


1
Expert's answer
2021-10-27T06:44:29-0400

V=πr2h=π·0.252·10≈1.9635cm3

Mass of Aluminum =1.9635×2.7=5.3036g

No. of aluminum atoms in the cylinder

=5.3036/4.48×10-23

=1.1838×1023 Al atoms